y = x+7 , x2 + y2 =29
Answer by Alan3354(69443)

y = x+7
x2 + y2 =29
---------------
Do you mean 2x + 2y = 29?
If so,
Sub for y in the 2nd eqn
2x + 2(x+7) = 29
2x + 2x + 14 = 29
4x = 15
x = 15/4, or 3.75
y = 10.75